# Quantum-Enhanced Power Flow and Optimal Power Flow based on Combinatorial Reformulation

### v2

@@ VI Results
The performance of AQPF and AQOPF is compared with NR provided by pandapower [23] in Section VI-A. These comparisons are performed based on the 118-bus test case [^5] https://icseg.iti.illinois.edu/ieee-118-bus-system/, which includes 118 buses, 173 branches, 13 transformers, 53 generators, and 54 loads. In Section VI-B, the performance of AQPF and AQOPF under ill-conditioned scenarios is evaluated for the 118-bus test case. In Section VI-C, a scalability analysis is presented for QIIO that demonstrates its ability to handle increasingly larger test cases.
@@
Figure 1: Performance comparison of the Newton-Raphson method (NR), Adiabatic Quantum Power Flow algorithm (AQPF), and partitioned AQPF for the 118-bus test case using Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO). The results are shown for (a) voltage magnitude $V_{i}\;\text{(p.u.)}$, (b) voltage phase angle $\delta_{i}\;\text{(degrees)}$, (c) net active power $P_{i}\;\text{(MW)}$, and (d) net reactive power $Q_{i}\;\text{(MVAR)}$.
@@
Figure 2: Performance comparison of the Newton-Raphson method (NR), Adiabatic Quantum Optimal Power Flow algorithm (AQOPF), and partitioned AQOPF for the 118-bus test case using Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO). The results are shown for (a) voltage magnitude $V_{i}\;\text{(p.u.)}$, (b) voltage phase angle $\delta_{i}\;\text{(degrees)}$, (c) net active power $P_{i}\;\text{(MW)}$, and (d) net reactive power $Q_{i}\;\text{(MVAR)}$.
VI-A Partitioned Formulation
@@
To assess the performance of the AQPF and AQOPF algorithms, their results, both in the full and partitioned formulations, are compared against NR for the 118-bus test case. The partitioned formulation is derived by excluding a subset of buses, denoted as $\mathbb{P}$, from the problem Hamiltonian in each iteration. Specifically, in each iteration, a set of 24 buses (approximately $20\%$ of the total buses) is randomly selected and removed from the computation. For PF analysis, this modification is applied in Algorithm 1 by excluding the subset $\mathbb{P}$ from the summation term in 15 at line 19, yielding:
[…]
@@
For OPF, the same subset $\mathbb{P}$ is excluded from the objective function in 29. This partitioned formulation allows for a reduced problem size per iteration.
VI-A1 Partitioned AQPF
@@
For PF analysis, the partitioned formulation in 35 results in a total of $7105$ binary variables (including $472$ base variables and auxiliary variables, which are introduced to reduce higher-order terms), reducing the variable count by $1775$ compared to the full AQPF algorithm. With a residual threshold of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, the mean square error (MSE) for the net active power $P_{i}$ between AQPF and NR is $4.28\times 10^{-4}\;\text{MW}^{2}$, while for partitioned AQPF, it is $8.06\times 10^{-4}\;\text{MW}^{2}$. Similarly, the MSE for the net reactive power $Q_{i}$ between AQPF and NR is $1.65\times 10^{-2}\;\text{MVAR}^{2}$, while for partitioned AQPF, it is $1.8\times 10^{-2}\;\text{MVAR}^{2}$. These results, illustrated in Fig. 1, indicate that AQPF closely matches NR, and the partitioned formulation does not significantly degrade solution quality. Further improvements can be achieved by reducing the residual threshold.
@@
VI-A2 Partitioned AQOPF
For OPF, the partitioned formulation reduces the number of binary variables (including $472$ base variables, auxiliary variables, which are introduced to reduce higher-order terms, and slack variables, which are introduced to integrate the constraints) to $8459$, which represents a reduction of $2189$ binary variables compared to the full formulation in 29. With a residual of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, the MSE for the net active power $P_{i}$ between AQOPF and NR is $8.89\times 10^{-4}\;\text{MW}^{2}$, while for partitioned AQOPF, it is $1.57\times 10^{-3}\;\text{MW}^{2}$. Similarly, the MSE for the net reactive power $Q_{i}$ between AQOPF and NR is $1.68\times 10^{-2}\;\text{MVAR}^{2}$, while for partitioned AQOPF, it is $1.69\times 10^{-2}\;\text{MVAR}^{2}$. The results are presented in Fig. 2.
@@ VI-A2 Partitioned AQOPF
VI-B Handling Ill-Conditioned Cases
TABLE II: Comparison of AQPF and AQOPF under well-conditioned and ill-conditioned scenarios for the 118-bus test case.
@@ VI-B Handling Ill-Conditioned Cases
Under these ill-conditioned scenarios, the NR solver in pandapower diverges for PF, and for OPF, numerical instability leads to infeasible solutions or failure to converge with optimization constraints remaining unmet. On the other hand, AQPF and AQOPF successfully solve both PF and OPF problems under the same ill-conditioned scenarios. The results are summarized in Table II.
The capability of AQPF and AQOPF in handling ill-conditioned cases is attributed to the enhanced numerical stability, that is, their combinatorial formulation avoids reliance on traditional linearization techniques, and hence mitigates issues caused by near-singular Jacobians. On the other hand, by iteratively refining increments/decrements, the algorithms ensure steady progress toward feasible solutions. The ability of Ising machines to handle a large number of binary decision variables also enables efficient searching of the solution space.
To further investigate the robustness of AQPF under ill-conditioned scenarios, the experiment is extended to include the partitioned AQPF formulation, as defined in 35, to solve PF analysis for the Ill-conditioned case 1. The results show that the partitioned AQPF formulation successfully solves the PF problem even under these challenging conditions. Specifically, the residual is $8.83\times 10^{-3}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, and the active and reactive power mismatches are $5.42\times 10^{-4}\;\text{MW}^{2}$ and $1.49\times 10^{-2}\;\text{MVAR}^{2}$, respectively. These results further underscore the capability of the partitioned AQPF approach to maintain computational efficiency without sacrificing solution accuracy, even in ill-conditioned scenarios.
VI-C Scalability
The scalability of the AQPF and AQOPF algorithms is demonstrated by their capacity to manage larger sets of binary variables using QIIO, which is confirmed through experiments conducted on test cases of different sizes. The results are summarized in Table III and Table IV, respectively; presenting the number of binary variables (-), the time needed to compile the QUBO formulation (s), the time elapsed per iteration (s), and the residual (${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$). Accordingly, QIIO is capable of handling the number of binary variables needed to solve PF up to the 1354-bus test case and the OPF problem up to the 300-bus test case. However, due to the high computational complexity associated with the larger test cases, only a limited number of iterations are performed for tehm with a residual threshold of $1\times 10^{0}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, serving as a proof of concept.
TABLE III: Scalability of AQPF for different test cases.
@@ VI-C Scalability
TABLE IV: Scalability of AQOPF for different test cases.
@@
VII Discussion
The following observations can be made:
@@ VII Discussion
  • D-Wave’s Advantage™  system (QA) and its hybrid quantum-classical solver (HA) are constrained by embedding challenges and available qubit count, which make them unsuitable for large test cases. While they successfully solve small-scale problems, their applicability to real-world power systems remains limited without improvements in embedding techniques, hardware connectivity, and qubit counts.
  • Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO) offers a remarkable capability of handling up to 100,000 binary variables, making it an excellent platform for realizing large-scale power system problems.
  • A more efficient update scheme can significantly enhance the performance of the proposed algorithms by minimizing computational overhead while maintaining accuracy. Future work, therefore, can improve the update scheme to reduce the number of iterations required for convergence.
@@
VIII Conclusion
This study introduces Adiabatic Quantum Power Flow (AQPF) and Adiabatic Quantum Optimal Power Flow (AQOPF) algorithms as novel approaches for, respectively, solving power flow (PF) and optimal power flow (OPF) problems using Ising machines. The feasibility of the proposed algorithms is demonstrated for large-scale power systems, and also their computational efficiency and scalability are highlighted. The findings confirm that the AQPF and AQOPF algorithms consistently achieve convergence across various test cases, including complex and ill-conditioned test cases, where conventional solvers often struggle. The results also suggest that Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO) offers a compelling alternative to classical solvers, particularly for problems with high-dimensional, combinatorial structures. This research, therefore, marks a significant step toward integrating Ising machines into power system applications but also paves the way for enhanced efficiency, scalability, and robustness in solving large-scale optimization problems.
@@ Acknowledgment
The authors gratefully acknowledge Heinz Wilkening and the service contract ”Quantum Computing for Load Flow” (contract number 690523) with the European Commission Directorate-General Joint Research Centre (EC DG JRC). Furthermore, the authors sincerely thank the Jülich Supercomputing Centre for allocating computing resources on the D-Wave Advantage™  System JUPSI via the Jülich UNified Infrastructure for Quantum computing (JUNIQ). This research was supported by the Center of Excellence RAISE, funded under the European Union’s Horizon 2020 Research and Innovation Framework Programme H2020-INFRAEDI-2019-1, grant agreement no. 951733. The authors also thank Fujitsu Technology Solutions for granting access to the QIIO software [^6] https://en-portal.research.global.fujitsu.com/kozuchi, with special appreciation to Markus Kirsch and Matthieu Parizy for their assistance and for providing customized extensions to the DADK Python package.

### v3

@@ 6 Results
The performance of AQPF and AQOPF is compared with NR provided by pandapower thurner2018pandapower in Section˜6.1. These comparisons are performed based on the 118-bus test case [^5] https://icseg.iti.illinois.edu/ieee-118-bus-system/, which includes 118 buses, 173 branches, 13 transformers, 53 generators, and 54 loads. In Section˜6.2, the performance of AQPF and AQOPF under ill-conditioned scenarios is evaluated for the 118-bus test case. In Section˜6.3, a scalability analysis is presented for QIIO that demonstrates its ability to handle increasingly larger test cases.
@@
Figure 1: Performance comparison of the NR method, AQPF, and partitioned AQPF for the 118-bus test case using QIIO. The results are shown for (a) voltage magnitude $V_{i}\;\text{(p.u.)}$, (b) voltage phase angle $\delta_{i}\;\text{(degrees)}$, (c) net active power $P_{i}\;\text{(MW)}$, and (d) net reactive power $Q_{i}\;\text{(MVAR)}$.
@@
Figure 2: Performance comparison of the NR method, AQOPF, and partitioned AQOPF for the 118-bus test case using QIIO. The results are shown for (a) voltage magnitude $V_{i}\;\text{(p.u.)}$, (b) voltage phase angle $\delta_{i}\;\text{(degrees)}$, (c) net active power $P_{i}\;\text{(MW)}$, and (d) net reactive power $Q_{i}\;\text{(MVAR)}$.
6.1 Partitioned Formulation
@@
To assess the performance of the AQPF and AQOPF algorithms, their results, both in the full and partitioned formulations, are compared against NR for the 118-bus test case. The partitioned formulation is derived by excluding a subset of buses, denoted as $\mathbb{P}$, from the problem Hamiltonian in each iteration. Specifically, in each iteration, a set of 24 buses (approximately $20\%$ of the total buses) is randomly selected and removed from the computation. For PF analysis, this modification is applied in Algorithm˜1 by excluding $\mathbb{P}$ from the summation term in ˜15 at line 19, yielding:
$H_{\text{part}}(\vec{x})=\sum_{i=1\atop i\notin\mathbb{P}}^{N=118}(P_{i}-P_{i}^{G}+P_{i}^{D})^{2}+(Q_{i}-Q_{i}^{G}+Q_{i}^{D})^{2}.$ (35)
@@
For OPF, the same subset $\mathbb{P}$ is excluded from the objective function in ˜29. This partitioned formulation reduces the problem size per iteration.
6.1.1 Partitioned AQPF
@@
For PF analysis, the partitioned formulation in ˜35 results in a total of $7105$ binary variables (including $472$ base variables and auxiliary variables, which are introduced to reduce higher-order terms), reducing the variable count by $1775$ compared to the full AQPF algorithm. With a residual threshold of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, the mean square error (MSE) for the net active power $P_{i}$ between AQPF and NR is $4.28\times 10^{-4}\;\text{MW}^{2}$, while for partitioned AQPF, it is $8.06\times 10^{-4}\;\text{MW}^{2}$. Similarly, the MSE for the net reactive power $Q_{i}$ between AQPF and NR is $1.65\times 10^{-2}\;\text{MVAR}^{2}$, while for partitioned AQPF, it is $1.8\times 10^{-2}\;\text{MVAR}^{2}$. These results, illustrated in Fig.˜1, indicate that AQPF closely matches NR, and the partitioned formulation does not significantly degrade solution quality. Further improvements can be achieved by reducing the residual threshold.
@@
6.1.2 Partitioned AQOPF
For OPF, the partitioned formulation reduces the number of binary variables (including $472$ base variables, auxiliary variables, which are introduced to reduce higher-order terms, and slack variables, which are introduced to integrate the constraints) to $8459$, which represents a reduction of $2189$ binary variables compared to the full formulation in ˜29. With a residual of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, the MSE for the net active power $P_{i}$ between AQOPF and NR is $8.89\times 10^{-4}\;\text{MW}^{2}$, while for partitioned AQOPF, it is $1.57\times 10^{-3}\;\text{MW}^{2}$. Similarly, the MSE for the net reactive power $Q_{i}$ between AQOPF and NR is $1.68\times 10^{-2}\;\text{MVAR}^{2}$, while for partitioned AQOPF, it is $1.69\times 10^{-2}\;\text{MVAR}^{2}$. The results are presented in Fig.˜2.
@@ 6.1.2 Partitioned AQOPF
6.2 Handling Ill-Conditioned Cases
Table 3: Comparison of AQPF and AQOPF under well-conditioned and ill-conditioned scenarios for the 118-bus test case.
@@ 6.2 Handling Ill-Conditioned Cases
Under these ill-conditioned scenarios, the NR solver in pandapower diverges for PF. For OPF, numerical instability leads to infeasible solutions or failure to converge with optimization constraints remaining unmet. On the other hand, AQPF and AQOPF successfully solve both PF and OPF problems under the same ill-conditioned scenarios. The results are summarized in Table˜3.
The capability of AQPF and AQOPF to handle ill-conditioned cases is attributed to their enhanced numerical stability; that is, their combinatorial formulation avoids reliance on traditional linearization techniques and hence mitigates issues caused by near-singular Jacobians. On the other hand, by iteratively refining increments/decrements, the algorithms ensure steady progress toward feasible solutions. The ability of Ising machines to handle a large number of binary decision variables also enables efficient searching of the solution space.
To further investigate the robustness of AQPF under ill-conditioned scenarios, the experiment is extended to include the partitioned AQPF formulation, as defined in ˜35, to solve PF analysis for the Ill-conditioned case 1. The results show that the partitioned AQPF formulation successfully solves the PF problem even under these challenging conditions. Specifically, the residual is $8.83\times 10^{-3}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, and the active and reactive power mismatches are $5.42\times 10^{-4}\;\text{MW}^{2}$ and $1.49\times 10^{-2}\;\text{MVAR}^{2}$, respectively. These results underscore the partitioned AQPF approach’s ability to maintain computational efficiency without sacrificing solution accuracy in ill-conditioned scenarios.
6.3 Scalability
The scalability of the AQPF and AQOPF algorithms is demonstrated by their ability to handle larger sets of binary variables using QIIO, as confirmed by experiments on test cases of varying sizes. The results are summarized in Table˜4 and Table˜5, respectively, presenting the number of binary variables (-), the time needed to compile the QUBO formulation (s), the time elapsed per iteration (s), and the residual (${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$). Accordingly, QIIO can handle the number of binary variables required to solve PF up to the 1354-bus test case and OPF up to the 300-bus test case. However, due to the high computational complexity associated with the larger test cases, only a limited number of iterations are performed for them with a residual threshold of $1\times 10^{0}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, serving as a proof of concept.
Table 4: Scalability of AQPF for different test cases.
@@ 6.3 Scalability
Table 5: Scalability of AQOPF for different test cases.
@@
7 Discussion
The results presented in this study should be interpreted within the broader context of algorithm development for emerging quantum computing paradigms. The intention of the proposed AQPF and AQOPF algorithms is not to compete with or replace established classical PF and OPF solvers, which have benefited from decades of research and remain highly effective for most practical applications. Instead, this study focuses on developing algorithms that could be useful as quantum hardware continues to mature. The proposed discrete combinatorial optimization approach provides an alternative mathematical representation of PF and OPF equations, enabling their implementation on Ising machines. At the same time, this representation leads to NP-hard optimization problems, and this complexity aligns with the capabilities targeted by quantum and digital annealers. Although current devices impose limitations on solution precision and problem scale, the experiments demonstrate that the methodology is computationally scalable and compatible with multiple annealing platforms. The following observations can be made:
@@ 7 Discussion
  • QA and HA are constrained by embedding challenges and available qubit count, making them unsuitable for large test cases. While they successfully solve small-scale problems, their applicability to real-world power systems remains limited without improvements in embedding techniques, hardware connectivity, and qubit counts.
  • QIIO offers a remarkable capability to handle up to 100,000 binary variables, making it an excellent platform for solving large-scale power system problems.
  • A more efficient update scheme can significantly enhance the performance of the proposed algorithms by minimizing computational overhead while maintaining accuracy. Future work can therefore improve the update scheme to reduce the number of iterations required for convergence.
@@
8 Conclusion
This study further investigates the PF and OPF problem reformulation as discrete combinatorial optimization problems that can be executed on Ising machines, including quantum and quantum-inspired hardware. Building upon the previously proposed Adiabatic Quantum Power Flow (AQPF) algorithm, this study introduces the Adiabatic Quantum Optimal Power Flow (AQOPF) algorithm. It evaluates the feasibility of both algorithms across a wide range of test cases. In addition, the computational efficiency and scalability of the proposed algorithms are highlighted. The findings confirm that the AQPF and AQOPF algorithms consistently converge across various test cases, including complex and ill-conditioned cases where conventional solvers often struggle. The results also suggest that QIIO offers a compelling alternative to classical solvers, particularly for problems with high-dimensional, combinatorial structures. This study, therefore, marks a significant step toward integrating Ising machines into power system applications and paves the way for enhanced efficiency, scalability, and robustness in solving large-scale optimization problems.
@@ Acknowledgment
This work was carried out within the DATALESs project (project no. 482.20.602), financially supported by the Netherlands Organization for Scientific Research (NWO) and the National Natural Science Foundation of China (NSFC). The authors gratefully acknowledge Heinz Wilkening and the service contract "Quantum Computing for Load Flow" (contract no. 690523) with the European Commission Directorate-General Joint Research Centre (EC DG JRC). Furthermore, the authors thank the Jülich Supercomputing Centre for allocating computing resources on the D-Wave Advantage™  System JUPSI via the Jülich UNified Infrastructure for Quantum computing (JUNIQ). This research was supported by the Center of Excellence RAISE, funded under the European Union’s Horizon 2020 Research and Innovation Framework Programme H2020-INFRAEDI-2019-1 (agreement no. 951733). The authors also thank Fujitsu Technology Solutions for granting access to the QIIO software [^6] https://en-portal.research.global.fujitsu.com/kozuchi, with special appreciation to Markus Kirsch and Matthieu Parizy for their assistance and for providing customized extensions to the DADK Python package.
[…]